(D3-2D2D')z=3x2y


1
Expert's answer
2021-10-22T01:04:35-0400

Given partial differential equation is,

"(D^3-2D^2D')z=3x^2y"


Let "D' = 1" then auxiliary equation,

"D^3-2D^2 = 0"

"m^3-2m^2 = 0 \implies m^2(m-2) = 0"

"m = 0, 0, 2"

Then

"C.F. , z = \phi(y)+x\phi(y)+\phi(y+2x)"


P.I. "\frac{1}{(D^3-2D^2D')}3x^2y = 3\frac{1}{(D^3-2D^2D')}x^2y = 3\frac{1}{D^3(1-\frac{2D'}{D})}x^2y"


Expanding biomomially,


"= \frac{3}{D^3}(1-\frac{2D'}{D})^{-1}x^2y = \frac{3}{D^3}(1+\frac{2D'}{D}+....)x^2y"


"= \frac{3}{D^3}[x^2y+\frac{2}{D}x^2]= \frac{3}{D^3}[x^2y+\frac{2x^3}{3}] = \frac{3x^2y}{D^3}+\frac{2x^3}{D^3}"


"= \frac{x^5y}{20}+\frac{x^6}{60}"



Complete solution will be,

"z = \phi(y)+x\phi(y)+\phi(y+2x) + \frac{x^5y}{20}+\frac{x^6}{60}"